# GSON: A Group-based Social Navigation Framework with Large Multimodal Model

### v2

@@ V-A Evaluation of LMM Social Group Prediction
To evaluate LMMs in predicting social groups from RGB images, we selected 100 scenarios from datasets [57, 58] and city walk videos, covering interactions such as queuing, walking, talking, and photographing. Human annotators provided group labels based on marked images, with the most agreed-upon annotations as ground truth. Example scenarios are shown in Fig. 6. On average, each scenario contains 6.0 pedestrians and 3.5 distinct social groups. We evaluated six LMMs for this task: four commercial models: Gemini 2.0 Flash, Gemini 2.5 Pro, Claude 3.7 Sonnet, GPT-4o, and two open-source models: Qwen2.5-VL (32B and 7B variants). Each scenario was tested 10 times with each model. The results were classified into four categories based on the predicted groupings: (1) Accurate: Group matches the ground truth (preferred), (2) Miss: Omits some individuals from the ground-truth group (acceptable), (3) Extra: Includes individuals not in the ground-truth group (not preferred), (4) Error: Both misses and includes incorrect individuals (not acceptable).  Fig. 7 shows the results for each model. All four LMMs achieved acceptable grouping rates above 67%, with the lowest from Qwen2.5-VL-7B (53.6% Accurate + 14.8% Miss) and the highest from GPT-4o at 86.6% (77.8% Accurate + 8.8% Miss). Given the complexity and diversity of the dataset, these results highlight the capability of Large Multimodal Models to identify social groups among pedestrians. Inference time varies across models: Gemini 2.5 Pro and Claude 3.7 Sonnet take around 8s, GPT-4o and Gemini 2.0 Flash about 4s, Qwen2.5-VL-32B around 2s, and Qwen2.5-VL-7B is the fastest, around 1s. In our pipeline, group detection doesn’t require high frequency, making GPT-4o and Qwen2.5-VL-32B well-suited for our system.
Figure 7: Grouping result of different LMMs in the dataset.
@@ V-B1 Experimental Setup
TABLE I: Social Evaluation Metrics of Planners
Metrics Explanation Calculation
[…]
Comfort Distance The average shortest distances between robot and group $\frac{1}{N}\sum_{i}d_{i}$, $d_{i}=\begin{cases}\min\|\mathbf{d}_{\mathrm{robot}}-\mathbf{d}_{\mathrm{convex}_{j}}\|&\text{robot not in convex of group}\\ -\max\|\mathbf{d}_{\mathrm{robot}}-\mathbf{d}_{\mathrm{convex}_{j}}\|&\text{robot in convex of group}\end{cases}$
@@ V-B2 Baseline and Evaluation Metrics
We compared GSON’s performance against several baselines, including classical planners like TEB and TEB enhanced with our group estimation, as well as recent open-source approaches: SACSON [23], an imitation learning-based method; GMPC [14], a group-based motion prediction controller; and RLCA [59], a distributed deep reinforcement learning method. Note that we excluded several LMM-based methods [27, 28, 25, 26] from the comparison because they are not open source and did not perform reliably in our pilot experiments with our own implementations. For evaluation, traditional navigation metrics included roughness, curvature, jerk, angular deviation from the planned path, and velocity [9, 60]. To assess social navigation, we introduced metrics to capture the robot’s impact on individuals and groups. Following prior guidelines [3], an individual was considered disturbed if the robot approached within 1.2 meters directly in front. For groups, a disturbance was defined by the robot entering the convex hull formed by group members. We measured both the duration of group disturbance and the ”comfort distance”—the average minimum distance between the robot and social groups.
@@ V-B3 Result and Analysis
The results are shown in Fig. 8. GSON performs comparably to other methods in traditional navigation metrics like Curvature, Jerk, and Roughness, achieving smooth and efficient motion with only a slight increase in jerk. In terms of social navigation, GSON clearly outperforms baselines. It records the lowest Time of Disturbing Individuals and Groups, showing minimal disruption to pedestrians and social groups. Notably, even compared to TEB+GROUP ESTIMATE, GSON maintains better group awareness. This is due to the short planning horizon of the TEB planner—when group members are spaced far apart, as in our scenarios, TEB+GROUP ESTIMATE can become stuck in local minima. GSON also achieves the highest Comfort Distance, maintaining more personal space from social groups—crucial in semantically rich environments. This demonstrates its strong respect for social norms during navigation. Overall, GSON balances traditional performance with superior social compliance, making it highly effective for real-world deployment in human-centric environments.
@@ VI Conclusion
In this work, we present a novel approach that integrates the visual reasoning capabilities of LMMs with a social structure-aware perception and planning system for mobile robots operating in human-centered environments. Our method utilizes the common sense reasoning ability of LMM by applying the visual prompting to predict the social relationships among pedestrians, which allows our planning system to generate socially aware behavior. Through extensive real-world navigation experiments, our approach demonstrated superior performance and robustness compared to baseline methods, illustrating its potential for improving social awareness and interaction in autonomous robot navigation. Future work could explore extending this framework to more dense social interactions, combining the system with compact state representation [61], and also distilling knowledge into smaller models for faster inference.

### v3

@@ V-A Evaluation of LMM Social Group Prediction
To evaluate LMMs in predicting social groups from RGB images, we selected 100 scenarios from datasets [53, 54] and city walk videos, covering interactions such as queuing, walking, talking, and photographing. Human annotators provided group labels based on marked images, with the most agreed-upon annotations as ground truth. Example scenarios are shown in Fig. 6. On average, each scenario contains 6.0 pedestrians and 3.5 distinct social groups. We evaluated six LMMs for this task: four commercial models: Gemini 2.0 Flash, Gemini 2.5 Pro, Claude 3.7 Sonnet, GPT-4o, and two open-source models: Qwen2.5-VL (32B and 7B variants). Each scenario was tested 10 times with each model. The results were classified into four categories based on the predicted groupings: (1) Accurate: Group matches the ground truth (preferred), (2) Miss: Omits some individuals from the ground-truth group (acceptable), (3) Extra: Includes individuals not in the ground-truth group (not preferred), (4) Error: Both misses and includes incorrect individuals (not acceptable).
Figure 7: Grouping results of LMMs and traditional clustering in the dataset
Fig. 7 shows the results for each model. All four LMMs achieved acceptable grouping rates above 67%, with the lowest from Qwen2.5-VL-7B (53.6% Accurate + 14.8% Miss) and the highest from GPT-4o at 86.6% (77.8% Accurate + 8.8% Miss). Given the complexity and diversity of the dataset, these results highlight the capability of Large Multimodal Models to identify social groups among pedestrians. Inference time varies across models: Gemini 2.5 Pro and Claude 3.7 Sonnet take around 8s, GPT-4o and Gemini 2.0 Flash about 4s, Qwen2.5-VL-32B around 2s, and Qwen2.5-VL-7B is the fastest, around 1s. In our pipeline, group detection doesn’t require high frequency, making GPT-4o and Qwen2.5-VL-32B well-suited for our system. For comparison, we evaluated two classical clustering baselines using DBSCAN on 2D and 3D coordinates. Both yielded significantly lower accuracy (10.8–14.8%) and higher extra grouping errors (41.9–46.5%), underscoring the limitations of proximity-based methods and the benefits of semantic reasoning in LMMs.
@@ V-B2 Baseline and Evaluation Metrics
We compared GSON’s performance against several baselines, including classical planners like TEB and TEB enhanced with our group estimation, as well as recent open-source approaches: SACSON [23], an imitation learning-based method; GMPC [14], a group-based motion prediction controller; and RLCA [55], a distributed deep reinforcement learning method. Note that we excluded several LMM-based methods [27, 28, 25, 26] from the comparison because they are not open source and did not perform reliably in our pilot experiments with our own implementations. For evaluation, traditional navigation metrics include roughness, curvature, jerk, and runtime [9, 56]. To assess social navigation, we introduced metrics to capture the robot’s impact on individuals and groups. Following prior guidelines [3], an individual was considered disturbed if the robot approached within 1.2 meters directly in front. For groups, a disturbance was defined by the robot entering the convex hull formed by group members. We measured both the duration of group disturbance and the “comfort distance”—the average minimum distance between the robot and social groups.
@@ V-B3 Result and Analysis
The results are shown in Fig. 8. GSON performs comparably to other methods in traditional navigation metrics like curvature, jerk, and roughness, achieving smooth and efficient motion with only a slight increase in jerk and runtime. Our speed is slightly lower than the comparison methods, mainly due to two reasons. First, calling the LMM adds at least 1 second of response time. Second, our method actively avoids groups, even if it means taking longer detours. However, the time is not significantly longer since the mid-level planning chooses the optimal path that does not interrupt the group. In terms of social navigation, GSON clearly outperforms baselines. It records the lowest Time of Disturbing Individuals and Groups, showing minimal disruption to pedestrians and social groups. Notably, even compared to TEB+GROUP ESTIMATE, GSON maintains better group awareness. This is due to the short planning horizon of the TEB planner—when group members are spaced far apart, as in our scenarios, TEB+GROUP ESTIMATE can become stuck in local minima. GSON also achieves the highest Comfort Distance, maintaining more personal space from social groups—crucial in semantically rich environments. This demonstrates its strong respect for social norms during navigation. Overall, GSON balances traditional performance with superior social compliance, making it highly effective for real-world deployment in human-centric environments.
@@ V-C Long-Range Social Navigation Demonstration
V-D Group-Aware Social Navigation in Dense Crowds: Limitations and Future Directions
To further investigate the ability of our system in high-density social environments and provide insights, we conducted additional experiments in three real-world scenarios: a university canteen during peak lunch hours, the entrance of a teaching building near an elevator, and the main campus gate. Each involved an average of over 13 pedestrians exhibiting dynamic, diverse social behaviors within the robot’s immediate workspace. While the robot successfully completed all navigation tasks, several limitations emerged under dense crowd conditions. We analyze these challenges and offer insights to guide future improvements.
LMM inference latency Latency from LMM-based group estimation—typically 1.5 to 2 seconds using Qwen2.5-VL-32B—sometimes reduced planner responsiveness to rapidly changing social contexts. To balance reasoning capability and speed, we plan to distill this ability into a smaller model, Qwen2.5-VL-3B, via fine-tuning.
Occlusions and partial observability Egocentric perception suffered from partial occlusions, where individuals blocked one another from the camera’s view, leading to incomplete or inaccurate group detection. In future work, we aim to incorporate predicted bird’s-eye views to infer social relationships beyond egocentric observations.
Complex and adaptive group behavior Currently, our framework outputs binary group labels (0 or 1), and the planner conservatively avoids all identified social groups. While sufficient for constructing conservative social spaces, this approach lacks flexibility to adapt to nuanced social contexts (e.g., traversing a queue may be more acceptable than interrupting a group chat). Incorporating finer-grained social understanding is a key next step. One promising direction is to integrate LMM-generated costs—based on different social interaction types—into the planning process.
@@ VI Conclusion
In this work, we present a novel approach that integrates the visual reasoning capabilities of LMMs with a social structure-aware perception and planning system for mobile robots operating in human-centered environments. Our method utilizes the common sense reasoning ability of LMM by applying the visual prompting to predict the social relationships among pedestrians, which allows our planning system to generate socially aware behavior. Through extensive real-world navigation experiments, our approach demonstrated superior performance and robustness compared to baseline methods, illustrating its potential for improving social awareness and interaction in autonomous robot navigation. Future work could explore extending this framework to more dense social interactions, combining compact state representation [57], and also distilling knowledge into smaller models for faster inference.